African Gold coffee sells for $12 per pound and Peruvian Dark sells for $8 per pound. How much of each do I need to use to make 50 pounds of a mixture that I will sell for $9?
Answer by stanbon(75887)

African Gold coffee sells for $12 per pound and Peruvian Dark sells for $8 per pound. How much of each do I need to use to make 50 pounds of a mixture that I will sell for $9?
----------------------
Equation:
value + value = value
12x + 8(50-x) = 9*50
-----
12 x + 400 - 8x = 450
----
4x = 50
x = 12.5 lbs (amt. of $12 coffee needed)
---
50-x = 37.5 lbs (amt of $8 coffee needed)
===================
Cheers,
Stan H.